Find regular singular point of (x^2+1)y"+y'+y=0


1
Expert's answer
2021-06-22T15:42:40-0400

"(x^2+1)y''+y'+y=0"


"a_2(x)=(x^2+1)=0"


"\implies" "x^2=-1"

Write the differential equation in standard form to get


"y"+{1\over (x^2+1)}y'+{1\over (x^2+1)}y=0"


"\implies p(x) ={1\over (x^2+1)}" and "q(x)={1\over (x^2+1)}"


Point "x^2=-1"

Both appear to the first power of p(x) and q(x) hence it gives to points which are both regular singular points.the points are

"x=\sqrt-1" and "x=-\sqrt-1"